OpenStudy (juan1857):

Use sum or difference identities to find the exact value of each trigonometric function: sinpie/2

OpenStudy (anonymous):

\[\sin(\frac{\pi}{2})\]?

OpenStudy (juan1857):

yeah

OpenStudy (anonymous):

you do not need an identity for this one, you should know this right off the bat

OpenStudy (anonymous):



OpenStudy (anonymous):

sine is the second coordinate

Nnesha (nnesha):

find two angles that add up to pi/2 (or difference = pi/2) \[\large\rm Sin(a-b)=\sin a \cos b - \cos a \sin b\] 2pi -3pi/2 = pi/2 \[\sin (2\pi -\frac{3\pi}{2}) = \sin 2\pi \cdot \cos \frac{3\pi}{2}- \cos2 \pi \cdot \sin \frac{3\pi}{2}\] solve

Nnesha (nnesha):

 hang this on wall you still have to use unit circle but i think ur teacher wants to see the work

OpenStudy (juan1857):

ok thanks

Nnesha (nnesha):

yw